OpenStudy (anonymous):

Given two component vectors, solve for the resultant vector. Use pythagorom theorm to find the hypotenuse and use inverse (arc) tangent to solve for the angle 1. x=-.75 y=-1.25 2. x=20000 y=14000

zepdrix (zepdrix):

\[\Large\rm \vec x=-.75,\qquad \vec y=-1.25\] \[\Large\rm R=\sqrt{(-.75)^2+(-1.25)^2}\approx1.46\] \[\Large\rm \theta=\arctan\left(\frac{-1.25}{-.75}\right)\approx 59^o\]This angle is in the first quadrant. We want the angle in the third quadrant (see how both our x and y components are negative?) \(\Large\rm 59+180=239^o\)

zepdrix (zepdrix):

What'd you end up with on the first one? :d anything like this?

OpenStudy (anonymous):

well i didn't really understand when u add or subtract 180?

OpenStudy (anonymous):

i understand whre u got r from and the 59

zepdrix (zepdrix):



zepdrix (zepdrix):

Arctangent, in order to be a function, has it's range restricted to these two quadrants. When we use the arctangent function, it will give us an angle between -90 degrees and 90 degrees. That's what's happening in our problem.